Grade 11Differential Calculus

if y=tan5x find dy/dx idsnoidsnvindiovnsodivnisnjvsdvojsnvkonskvnksnvkosnkvnskovnsknvksnvknskvnksnvkdnvkondfkvndkovnkodnvdnv

RAKESH
8 Years agoGrade 11

3 Answers

sonika p
8 Years ago
according to chain rule, we first differentiate the exponential term according to
d/dt (zt) = tzt-1,
and then multiply by the differentiation of tanx which is sec2x. 
dy/dx = (5tan4x) * (sec2x) = 5tan4x sec2x
 
Chandan
8 Years ago
dy/dx=d(tan5x)/dx
=d(tan5x)/d(tanx) * d(tanx)/dx multiplying and dividing by d(tanx)
=5tan4xsec2x.
hope it helps 
please approve if it does
Sanjana
8 Years ago
Substitute u = g(x). This gives usy = f(u)Next we need to use a formula that is known as the Chain Rule.dy/dx = dy/du × du/dx Dy/dx=5tan^4 x .d (tan x)/dx =5sec^2 x tan^4 x